OpenStudy (anonymous):

Please Help! I Need Explanation & Help On How To Do This!!!! Medal and Fan?!?!

OpenStudy (anonymous):

√(8x+1)=5

OpenStudy (anonymous):

i need to solve for x and identify if it is an extraneous solution.

OpenStudy (anonymous):

Is it \(\huge \sqrt{8x+1}=5\) ?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

\(\huge (\sqrt{8x+1})^2=(5)^2\)

OpenStudy (anonymous):

where did you get the power of 2?

OpenStudy (anonymous):

So you square both sides to remove the radical sign.

OpenStudy (anonymous):

ohhhh nvm wait

OpenStudy (anonymous):

i understand why lol

OpenStudy (anonymous):

It would now be: \(\huge 8x+1=25\)

OpenStudy (anonymous):

so x equals 3??

OpenStudy (anonymous):

Yes! To check if it's an extraneous solution, plug in \(\huge x=3\) to the equation.

OpenStudy (anonymous):

\(\huge \sqrt{8(3)+1}=5; x=3 \)

OpenStudy (anonymous):

whats the difference between non extraneous and extraneous?

OpenStudy (anonymous):

extraneous: not a real solution non extraneous: real solution / the x/root is true

OpenStudy (anonymous):

\(\huge \sqrt{24+1}=5\) \(\huge \sqrt{25}=5\) Is this true?

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

ok thank you so much!!!